Hello. Just want to make sure I have this correct.
Solve the following absolute value equation.
|x-3| = 100
x-3+3=-100+3
x=-97
x-3+3=100+3
x=103
If someone can check this for me, I'd appreciate it. Thanks!

Answer by stanbon(75887)

|x-3| = 100
x-3=100 or -(x-3)=100
x=103 or x-3=-100
x=103 or x=-97
Cheers,
Stan H.